I need to know STEP BY STEP how to work out this problem. 3(2x-5)=4(x=3) and how to get the answer.. please i really really need help!! Thanks so much!
Answer by checkley71(8403)

3(2X-5)=4(X+3) I'VE ASSUMED YOU MENT TO HAVE A + BETWEEN THE X&3?????
6X-10=4X+12 OR 6X-4X=12+10 OR 2X=22 OR X=11